the complete combustion of candle wax (C31H64) in oxygen produced co2 and liquid water. calculate the work associated with this combustion reaction at 298k in kj/mol
Answer #1

Balanced equation : 2 C31H64(S) + 94 O2(g) ----> 62 CO2(g) + 64 H2O(l)

Dn = np-nR = 62- 94 = 32

Work done(w) = -DnRT

             = -32*8.314*298

             = -79.28 kj

per mole of Wax(C31H64) Work done(w) = -79.28/2 = - 39.64 kj

Answer: w = -39.64 kj/mol